OpenStudy (anonymous):

find the indicatedputs for f(x) 2x2-5x f(0)= f(-1)= f(2)=

OpenStudy (anonymous):

whats the problem here?

OpenStudy (anonymous):

Plugging in for x=0, we have: 2(0)^2 - 5(0) = 0 f(-1) = 2(-1)^2-5(-1) = 2+5 = 7 f(2) = 2(2)^2-5(2) = 8-10 = -2 Hope this helps! http://www.tutorsean.net Follow me on Facebook: Tutor Sean